# Volumetric Functional Maps

### v1

@@ Abstract.
The computation of volumetric correspondences between 3D shapes has great potential for medical and industrial applications. In this work, we pave the way for spectral volume mapping, extending for the first time the functional maps framework from the surface setting to the volumetric domain. We show that the eigenfunctions of the volumetric Laplace operator define a functional space that is suitable for high-quality signal transfer. We also experiment with various techniques that edit this functional space, porting them from the surface to the volume setting. We validate our method on novel volumetric datasets and on tetrahedralizations of well established surface datasets, also showcasing practical applications involving both discrete and continuous signal mapping, for segmentation transfer, mesh connectivity transfer and solid texturing. Last but not least, we show that considering the volumetric spectrum greatly improves the accuracy for classical shape matching tasks among surfaces, consistently outperforming existing surface-only spectral methods.
Functional maps; volume mapping; shape matching; shape correspondence
submissionid: 598ccs: Computing methodologies Shape analysisccs: Computing methodologies Volumetric modelsccs: Computing methodologies Computer graphics
@@ 1. Introduction
The computation of volumetric correspondences between solid objects is becoming more and more important for applications. To make some concrete examples, volume mappings enable non-invasive medical diagnosis, by warping the digital copy of a human organ into its canonical position without extracting it from the patient’s body (Abulnaga et al., 2021). They are also important for hexmeshing (Pietroni et al., 2022), statistical shape analysis (Huang et al., 2019) and for Industry 4.0, where physical simulations computed on a predictive digital twin must be aligned with the other digital views of the same real environment for optimized decision making and failure avoidance (Rasheed et al., 2020).
Maps between 3D objects are often represented in a computer either by exploiting one-to-one per vertex correspondences between two simplicial meshes with same connectivity, defining a so called piece-wise map (Section 2.1), or putting into correspondence two sets of basis functions that allow to express the same signal in two alternative domains, as done by spectral approaches for surface meshes (Ovsjanikov et al., 2012). Both piece-wise and spectral approaches have been extensively studied in the surface setting and can nowadays be considered rather mature, as proved by the numerous academic and commercial tools that incorporate these technologies, such as Blender, Maya and 3DS Max. Conversely, the equivalent task of computing volumetric correspondences between two solid objects largely remains a scientific open problem (Pietroni et al., 2022, §8.2).
@@
In this article, we propose this extension for the first time and show that, conversely from the piece-wise linear setting, functional mapping naturally scales to volumes with minimal effort, providing a theoretically sound and practically useful platform for the computation of volumetric correspondences between solid shapes.
Using the spectrum of the volumetric Laplacian we define a functional space that is completely agnostic w.r.t. the underlying discrete mesh and we use it to create correspondences between solid meshes with different density and connectivity. Interestingly, in the paper we show that many recent developments used to improve efficiency by reducing the size of the functional space without sacrificing accuracy (Melzi et al., 2019b; Maggioli et al., 2021) can also be exploited in the volume setting, permitting to find an optimal balance between map quality and performances.
@@
All in all, this article opens for the first time to the computation of volumetric correspondences with spectral approaches. To this end, we believe our contributions will foster new research in this direction, positioning volume spectral mapping as a flexible and practical alternative to existing piece-wise linear approaches.
@@ 2.1. Piece-wise Maps
A less robust but more practical alternative to the map composition pipeline consists in initializing a non injective volume map directly, with Tutte 3D or some other non robust method, and then restore the correct orientation for elements that flip or vanish through the map via vertex relocation. In the volume setting orientation (injectivity) constraints translate to sign constraints on the determinant of the $3\times 3$ Jacobian matrix associated to each per element transformation, leading to an optimization problem subject to cubic constraints which is numerically intractable. Most of the research in the field focuses on how to formulate (and relax) the orientation constraints to make the problem tractable. Removing inverted elements from a given mesh is also a well known problem in computer simulation, where it is often referred to as mesh untangling (Knupp, 2001; Pietroni et al., 2022). Note that in the general case mesh untangling is an ill-posed problem, because it may not admit a valid solution for a fixed mesh connectivity. Additional degrees of freedom must be inserted via mesh refinement to open the feasible region (see (Livesu, 2020b, Fig.1) for a 2D example). Existing volume approaches employ minimization of flip-preventing energies (Du et al., 2020; Garanzha et al., 2021; Abulnaga et al., 2023), operate on convex subsets of the feasible region (Kovalsky et al., 2014) or project onto the space of bounded distortion mappings (Aigerman and Lipman, 2013). Due to the non linearity of the problem and the non existence of a feasible starting solution, none of these methods guarantees the correctness of the result. Multiple failure cases have been reported by testing the most prominent existing algorithms on available datasets (Nigolian et al., 2023, §6.3).
@@ 3. Background and Notation
We discretize volumes as tetrahedral meshes $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$, where:
  • $V_{\mathcal{M}}\subset\mathbb{R}^{3}$ is a set of vertices in 3D space;
  • $T_{\mathcal{M}}\subset V_{\mathcal{M}}^{4}$ is a set of tetrahedra connecting vertices in $V_{\mathcal{M}}$.
@@
Given a tetrahedral (or surface) mesh $\mathcal{M}$, and given two vertices $v_{i},v_{j}\in\mathcal{M}$, we refer to the geodesic distance between $v_{i}$ and $v_{j}$ over $\mathcal{M}$ as $d_{\mathcal{M}}(v_{i},v_{j})$. There exist several algorithms to estimate such a distance on discrete meshes (Crane et al., 2020). For the sake of simplicity, in the remainder of the paper we will always measure $d_{\mathcal{M}}(v_{i},v_{j})$ using Dijkstra’s algoritm (Dijkstra, 1959) on the edge connectivity of $\mathcal{M}$.
@@
Given two sets of (orthogonal) basis functions $\mathbf{\Phi}_{\mathcal{M}},\mathbf{\Phi}_{\mathcal{N}}$, respectively for the functional space $\mathcal{F}(\mathcal{M}),\mathcal{F}(\mathcal{N})$, and a function $f:V_{\mathcal{N}}\to\mathbb{R}$, the function $f$ can be represented as a linear combination of the basis functions $\mathbf{\Phi}_{\mathcal{N}}$ as
@@
where $\phi_{\mathcal{N}}^{(i)}$ is the $i$-th basis function in $\mathbf{\Phi}_{\mathcal{N}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{N}}$ denotes the inner product between functions over $\mathcal{N}$. The coefficients $\alpha_{i}$ thus denotes the projection of $f$ onto the basis functions $\phi_{\mathcal{N}}^{(i)}$. A usual choice for the bases $\mathbf{\Phi}_{\mathcal{M}},\mathbf{\Phi}_{\mathcal{N}}$ are the eigenfunctions of the Laplace-Beltrami operator on $\mathcal{M}$ and $\mathcal{N}$, respectively.
@@
(4) $T_{\pi}(f)\approx\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}\bm{f}\,,$
@@
Notably, $T_{\pi}(\mathbf{\Phi}_{\mathcal{N}})\approx\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}$, and thus by knowing the matrix $\mathbf{C}$ it is possible to extract the unknown correspondence $\pi$ via a nearest-neighbor search $\mathrm{NN}(\mathbf{\Phi}_{\mathcal{N}},\mathbf{\Phi}_{\mathcal{M}}\mathbf{C})$.
@@ 4. Method
The functional maps framework is not bound to the domain of 2-dimensional surfaces. Indeed, the Laplace-Beltrami Operator (LBO) can be defined on Riemannian manifolds of arbitrary dimension, and it always admits a spectral decomposition. In this work, we discretize the LBO of a tetrahedral mesh $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ using the $n$-dimensional cotangent formula (Liao et al., 2009; Crane, 2019), which defines the discrete LBO as a pair of matrices $\mathbf{S}$ and $\mathbf{W}$, respectively denoted as stiffness and mass matrices. The spectral decomposition of the LBO can therefore be obtained by solving the generalized eigenproblem $\mathbf{S}\mathbf{\Phi}=\mathbf{W}\mathbf{\Phi}\mathbf{\Lambda}$. However, most of the properties related to the LBO eigenbasis that are regularly exploited by functional maps algorithms cannot be taken for granted due to the structural differences between tetrahedral and triangular meshes.
@@ 4.1. Functional Maps for Tetrahedral Meshes
Most successful algorithms for the estimation of the functional map between shapes rely on the property of intrinsic structures and processes to behave similarly on similar surfaces. For instance, the invariance of the LBO basis and the heat diffusion under isometric and quasi-isometric deformations plays a key role in stable functional map estimation methods (Nogneng and Ovsjanikov, 2017; Nogneng et al., 2018; Ovsjanikov et al., 2016).
@@
To ensure that the LBO has a consistent behavior across isometric or similar volumetric shapes, we compute the spectral decomposition of the LBO on 40 pairs of shapes with identical connectivity released by Su et al. (2019) and compare both the eigenvalues and the eigenfunctions of the LBO. Our results are summarized in Figure 2. The first two rows of the figure show that two humanoid shapes share the same LBO eigenfunctions (up to sign flips), even if they are not isometric and not representing the same subject: the overall similarity is sufficient to generate the same low-frequency spectrum. To compare the eigenvalues, we first compute the vectors $\bm{\lambda}_{\mathcal{M}},\bm{\lambda}_{\mathcal{N}}$ of the first 100 LBO eigenvalues on two tetrahedral meshes, and then evaluate the difference $|\bm{\lambda}_{\mathcal{M}}-\bm{\lambda}_{\mathcal{N}}|$ Since the absolute difference between LBO eigenvalues is notoriously not informative, we compute both the relative difference and the offset difference proposed by Moschella et al. (2022). For each pair of tetrahedral meshes, we also extract the triangular meshes representing their boundary surface and compute the same quantities. The error distributions are shown at the bottom of Figure 2, where we can see that the spectral similarity between pairs of tetrahedral and surface meshes is comparable.
@@
(5) $\|\mathbf{C}\|_{O}=\frac{\|\mathbf{C}^{\top}\mathbf{C}-\mathbf{I}\|}{\|\mathbf{I}\|}\,,\qquad\|\mathbf{C}\|_{L}=\frac{\|\mathbf{C}\mathbf{\Lambda}_{\mathcal{M}}-\mathbf{\Lambda}_{\mathcal{N}}\mathbf{C}\|}{\|\mathbf{C}\|}\,,$
@@ 4.2. Spectral Connectivity Transfer
Given the surface $\partial\mathcal{N}$ and its induced tetrahedral mesh $\mathcal{N}$, we exploit the known correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$ to induce a full correspondence $\pi^{\prime}:\mathcal{M}\to\mathcal{N}$. Let $\mathbf{\Phi}_{\mathcal{M}},\mathbf{\Phi}_{\mathcal{N}}$ be the LBO eigenbases of the volumes $\mathcal{M}$ and $\mathcal{N}$, respectively, and let us denote with $\mathbf{\Phi}_{\partial\mathcal{M}}=(\mathbf{\Phi}_{\mathcal{M}})_{\partial},\mathbf{\Phi}_{\partial\mathcal{N}}=(\mathbf{\Phi}_{\mathcal{N}})_{\partial}$ their restrictions to the surfaces $\partial\mathcal{M}$ and $\partial\mathcal{N}$, respectively. We stress that $\mathbf{\Phi}_{\partial\mathcal{M}}$ is not the LBO eigenbasis of $\partial\mathcal{M}$. We note that the functional map $\mathbf{C}$ induced by $\pi^{\prime}$ must be such that
(6) $\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}=T_{\pi^{\prime}}(\mathbf{\Phi}_{\mathcal{N}})\,.$
We consider that the rows of $\mathbf{\Phi}_{\partial\mathcal{M}}$ (resp. $\mathbf{\Phi}_{\partial\mathcal{N}}$) are a subset of the rows of $\mathbf{\Phi}_{\mathcal{M}}$ (resp. $\mathbf{\Phi}_{\mathcal{N}})$. Recalling that $\pi$ is the restriction of $\pi^{\prime}$, we get
(7) $\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{C}=(\mathbf{\Phi}_{\mathcal{M}})_{\partial}\mathbf{C}=(\mathbf{\Phi}_{\mathcal{M}}\mathbf{C})_{\partial}=(T_{\pi^{\prime}}(\mathbf{\Phi}_{\mathcal{N}}))_{\partial}=T_{\pi}(\mathbf{\Phi}_{\partial\mathcal{N}})\,.$
The size of the basis (i.e., the number of columns of $\mathbf{\Phi}_{\partial\mathcal{M}}$) is generally smaller than the number of vertices (i.e., the number of rows), therefore $\mathbf{\Phi}_{\partial\mathcal{M}}$ admits left inverse and the volumetric functional map can be approximated via the known surface correspondence as $\mathbf{C}\approx\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(\mathbf{\Phi}_{\partial\mathcal{N}})$.
@@
(8) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}x_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}y_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}z_{\mathcal{N}}\,.\end{gathered}$

### v2

@@ Abstract.
The computation of volumetric correspondences between 3D shapes is a prominent tool for medical and industrial applications. In this work, we pave the way for spectral volume mapping, extending for the first time the functional maps framework from the surface to the volumetric setting. We show that the eigenfunctions of the volumetric Laplace operator define a functional space that is suitable for high-quality signal transfer. We also experiment with various techniques that edit this functional space, porting them to volume domains. We validate our method on novel volumetric datasets and on tetrahedralizations of well established surface datasets, also showcasing practical applications involving both discrete and continuous signal mapping, for segmentation transfer, mesh connectivity transfer and solid texturing. Last but not least, we show that considering the volumetric spectrum greatly improves the accuracy for classical shape matching tasks among surfaces, consistently outperforming existing surface-only spectral methods.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: TOGjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 1. Introduction
The computation of correspondences between volumetric objects plays a central role for many critical applications. Volume mappings enable non-invasive medical diagnosis, by warping the digital copy of a human organ into its canonical position without extracting it from the patient’s body (Abulnaga et al., 2021). They are also important for hexmeshing (Pietroni et al., 2022), statistical shape analysis (Huang et al., 2019) and for Industry 4.0, where physical simulations computed on a predictive digital twin must be aligned with the other digital views of the same real environment for optimized decision making and failure avoidance (Rasheed et al., 2020).
Maps between 3D objects are often represented in a computer either by exploiting one-to-one per vertex correspondences between two simplicial meshes with same connectivity, defining a so called piece-wise map (Section 2.1), or putting into correspondence two sets of basis functions that allow to express the same signal in two alternative domains, as done by spectral approaches for surface meshes (Ovsjanikov et al., 2012). Both piece-wise and spectral approaches have been extensively studied in the surface setting and can nowadays be considered rather mature, as proved by the numerous academic and commercial tools that incorporate these technologies, such as Blender, Maya and 3DS Max. Conversely, the equivalent task of computing volumetric correspondences between solid objects largely remains a scientific open problem (Pietroni et al., 2022; Nigolian et al., 2023, §8.2).
@@
In this article, for the first time, we propose this extension and show that, conversely from the piece-wise linear setting, functional mapping naturally scales to volumes with minimal effort, providing a theoretically sound and practically useful platform for the computation of volumetric correspondences between solid shapes.
Using the spectrum of the volumetric Laplacian, we define a functional space that is completely agnostic with respect to the underlying discrete mesh, and we use it to create correspondences between solid meshes with different density and connectivity. Interestingly, in the paper we show that many recent developments used to improve efficiency by reducing the size of the functional space without sacrificing accuracy (Melzi et al., 2019b; Maggioli et al., 2021) can also be exploited in the volume setting, permitting the finding of an optimal balance between map quality and performance.
@@
All in all, this article opens for the first time to the computation of volumetric correspondences with spectral approaches. To this end, we believe our contributions will foster new research in this direction, positioning volume spectral mapping as a flexible and practical alternative to existing piece-wise linear approaches. To facilitate this endeavor, we release our reference code at anonymous.
@@ 2.1. Piece-wise Maps
A less robust but more practical alternative to the map composition pipeline consists in initializing a non injective volume map directly, with Tutte 3D or some other non robust method, and then restore the correct orientation for elements that flip or vanish through the map via vertex relocation. In the volume setting, orientation (injectivity) constraints translate to sign constraints on the determinant of the $3\times 3$ Jacobian matrix associated to each per element transformation, leading to an optimization problem subject to cubic constraints which is numerically intractable. Most of the research in the field focuses on how to formulate (and relax) the orientation constraints to make the problem tractable. Removing inverted elements from a given mesh is also a well known problem in computer simulation, where it is often referred to as mesh untangling (Knupp, 2001; Pietroni et al., 2022). Note that in the general case mesh untangling is an ill-posed problem, because it may not admit a valid solution for a fixed mesh connectivity. Additional degrees of freedom must be inserted via mesh refinement to open the feasible region (see (Livesu, 2020b, Fig.1) for a 2D example). Existing volume approaches employ minimization of flip-preventing energies (Du et al., 2020; Garanzha et al., 2021; Abulnaga et al., 2023), operate on convex subsets of the feasible region (Kovalsky et al., 2014) or project onto the space of bounded distortion mappings (Aigerman and Lipman, 2013). Due to the non linearity of the problem and the non existence of a feasible starting solution, none of these methods guarantees the correctness of the result. Multiple failure cases have been reported by testing the most prominent existing algorithms on available datasets (Nigolian et al., 2023, §6.3). In Section 5.2 we show that bootstrapping this pipeline with our functional mapping dramatically increases the success rate of state of the art volume untangling algorithms.
@@ 3. Background and Notation
We discretize volumes as tetrahedral meshes $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$, where: $V_{\mathcal{M}}\subset\mathbb{R}^{3}$is a set of vertices in 3D space; $T_{\mathcal{M}}\subset V_{\mathcal{M}}^{4}$is a set of tetrahedra connecting the vertices in $V_{\mathcal{M}}$.
@@
Given a tetrahedral (or triangular) mesh $\mathcal{M}$, and given two vertices $v_{i},v_{j}\in\mathcal{M}$, we refer to the geodesic distance between $v_{i}$ and $v_{j}$ over $\mathcal{M}$ as $d_{\mathcal{M}}(v_{i},v_{j})$. There exist several algorithms to estimate such a distance on discrete meshes (Crane et al., 2020). For the sake of simplicity, in the remainder of the paper we will always measure $d_{\mathcal{M}}(v_{i},v_{j})$ using Dijkstra’s algoritm (1959) on the edge connectivity of $\mathcal{M}$.
@@
Given two sets of (orthogonal) basis functions $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$, respectively for the functional space $\mathcal{F}(\mathcal{M}),\mathcal{F}(\mathcal{N})$, and a function $f:V_{\mathcal{N}}\to\mathbb{R}$, the function $f$ can be represented as a linear combination of the basis functions $\Phi_{\mathcal{N}}$ as
@@
where $\phi_{\mathcal{N}}^{(i)}$ is the $i$-th basis function in $\Phi_{\mathcal{N}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{N}}$ denotes the inner product between functions over $\mathcal{N}$. The coefficients $\alpha_{i}$ thus denotes the projection of $f$ onto the basis functions $\phi_{\mathcal{N}}^{(i)}$. A usual choice for the bases $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ are the eigenfunctions of the Laplace-Beltrami operator on $\mathcal{M}$ and $\mathcal{N}$, respectively.
@@
(4) $T_{\pi}(f)\approx\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}\bm{f}\,,$
@@
Notably, $T_{\pi}(\Phi_{\mathcal{N}})\approx\Phi_{\mathcal{M}}\mathbf{C}$, and thus by knowing the matrix $\mathbf{C}$ it is possible to extract the unknown correspondence $\pi$ via a nearest-neighbor search $\mathrm{NN}(\Phi_{\mathcal{N}},\Phi_{\mathcal{M}}\mathbf{C})$.
@@ 4. Method
The functional maps framework is not bound to the domain of 2-dimensional surfaces. Indeed, the Laplace-Beltrami Operator (LBO) can be defined on Riemannian manifolds of arbitrary dimension and it always admits a spectral decomposition. We discretize the LBO of a tetrahedral mesh $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ using the $n$-dimensional cotangent formula (Liao et al., 2009; Crane, 2019) which defines the discrete LBO as a pair of matrices $\mathbf{S}$ and $\mathbf{W}$, respectively denoted as stiffness and mass matrices. Its spectral decomposition can be obtained by solving the generalized eigenproblem $\mathbf{S}\Phi=\mathbf{W}\Phi\Lambda$.
We provide in Appendix A the implementation details, as well as a brief discussion about the choice of the LBO and its discretization. However, we highlight that, since most applications that we address involve analyzing information at the surface (which is the boundary of a tetrahedral mesh), we are forced to use a discretization that imposes Neumann’s boundary conditions to avoid zero-valued eigenfunctions on the surface.
@@ 4.1. Functional Maps for Tetrahedral Meshes
Most successful algorithms for the estimation of the functional map between shapes rely on the property of intrinsic structures and processes to behave similarly on similar surfaces. For instance, the invariance of the LBO basis and the heat diffusion under isometric and quasi-isometric deformations plays a key role in stable functional map estimation methods (Nogneng and Ovsjanikov, 2017; Nogneng et al., 2018; Ovsjanikov et al., 2016). We provide a discussion on isometric deformations between volumes in Appendix C.
@@
To ensure that the LBO has a consistent behavior across isometric or similar volumetric shapes, we compute the spectral decomposition of the LBO on 40 pairs of shapes with identical connectivity released by Su et al. (2019) and compare both the eigenvalues and the eigenfunctions of the LBO. Our results are summarized in Figure 2. The first two rows of the figure show that two humanoid shapes share the same LBO eigenfunctions (up to sign flips), even if they are not isometric and not representing the same subject: the overall similarity is sufficient to generate the same low-frequency spectrum. To compare the eigenvalues, we first compute the vectors $\bm{\lambda}_{\mathcal{M}},\bm{\lambda}_{\mathcal{N}}$ of the first 100 LBO eigenvalues on two tetrahedral meshes. Since the absolute difference between LBO eigenvalues is notoriously not informative, we rely on both the relative difference and the offset difference proposed by Moschella et al. (2022). For a deeper discussion about the error measures, we refer to Appendix B.
For each pair of tetrahedral meshes, we also extract the triangular meshes representing their boundary surface and compute the same quantities. For a more meaningful comparison, all the tetrahedral meshes are rescaled to have unitary volume, and all the surface meshes are rescaled to have unitary surface area. The error distributions are shown at the bottom of Figure 2, where we can see that the spectral similarity between pairs of tetrahedral and surface meshes is comparable.
@@
(5) $\|\mathbf{C}\|_{O}=\frac{\|\mathbf{C}^{\top}\mathbf{C}-\mathbf{I}\|}{\|\mathbf{I}\|}\,,\qquad\|\mathbf{C}\|_{L}=\frac{\|\mathbf{C}\Lambda_{\mathcal{M}}-\Lambda_{\mathcal{N}}\mathbf{C}\|}{\|\mathbf{C}\Lambda_{\mathcal{M}}\|}\,,$
@@ 4.2. Spectral Connectivity Transfer
Given the surface $\partial\mathcal{N}$ and its induced tetrahedral mesh $\mathcal{N}$, we exploit the known correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$ to induce a full correspondence $\pi^{\prime}:\mathcal{M}\to\mathcal{N}$. Let $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ be the LBO eigenbases of the volumes $\mathcal{M}$ and $\mathcal{N}$, respectively, and let us denote with $\Phi_{\partial\mathcal{M}}=(\Phi_{\mathcal{M}})_{\partial},\Phi_{\partial\mathcal{N}}=(\Phi_{\mathcal{N}})_{\partial}$ their restrictions to the surfaces $\partial\mathcal{M}$ and $\partial\mathcal{N}$, respectively. We stress that $\Phi_{\partial\mathcal{M}}$ is not the LBO eigenbasis of $\partial\mathcal{M}$. We note that the functional map $\mathbf{C}$ induced by $\pi^{\prime}$ must be such that
(6) $\Phi_{\mathcal{M}}\mathbf{C}=T_{\pi^{\prime}}(\Phi_{\mathcal{N}})\,.$
We consider that the rows of $\Phi_{\partial\mathcal{M}}$ (resp. $\Phi_{\partial\mathcal{N}}$) are a subset of the rows of $\Phi_{\mathcal{M}}$ (resp. $\Phi_{\mathcal{N}})$. Recalling that $\pi$ is the restriction of $\pi^{\prime}$, we get
(7) $\Phi_{\partial\mathcal{M}}\mathbf{C}=(\Phi_{\mathcal{M}})_{\partial}\mathbf{C}=(\Phi_{\mathcal{M}}\mathbf{C})_{\partial}=(T_{\pi^{\prime}}(\Phi_{\mathcal{N}}))_{\partial}=T_{\pi}(\Phi_{\partial\mathcal{N}})\,.$
In general, the size of the basis (i.e., the number of columns of $\Phi_{\partial\mathcal{M}}$) is significantly smaller than the number of vertices (i.e., the number of rows), even if we only consider the vertices at the surface. Intuitively, this suggests that the restricted eigenfunctions $\Phi_{\partial\mathcal{M}}$ at the surface are linearly independent. Indeed, classical results from control theory support this intuition, proving that the boundary traces (i.e., the restriction to the boundary) of the eigenfunctions of any elliptic operator are linearly independent in a system with Neumann boundary conditions (Lasiecka and Triggiani, 1983; Triggiani, 2008) (see Appendix D for a deeper analysis of the descriptive power of the boundary traces). Therefore, $\Phi_{\partial\mathcal{M}}$ admits left inverse and the volumetric functional map can be approximated via the known surface correspondence as $\mathbf{C}\approx\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(\Phi_{\partial\mathcal{N}})$.
@@
(8) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}x_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}y_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}z_{\mathcal{N}}\,.\end{gathered}$
@@
Figure 4. A visual representation of the pipeline for extrapolating the interior coordinates from the surface correspondence. Given the surface correspondence $\pi$ (first row), we approximate the spectral embedding of the surface coordinates of $\partial\mathcal{N}$ using the boundary restriction of the eigenfunctions of $\mathcal{M}$ (second row). Then, we reconstruct the interior coordinates from the spectral embedding using the eigenfunctions on the entire volume (third row), and we use the reconstructed coordinates to transfer the interior connectivity (fourth row).